Represent -1/4,1/5 and 7/5 on a number line
Answer by CubeyThePenguin(3113)

-1/4 = -0.25
1/5 = 0.2
7/5 = 1.4

-1..........0..........1..........2
-----------------------------
..........|.....|.........|